OpenStudy (anonymous):

Prove cscØ/cotØ - cotØ/cscØ=sin^2Ø/cosØ

OpenStudy (anonymous):

probably no trig at all involved with this, just algebra if you put \(\cos(\theta)=a\) and \(\sin(\theta)=b\) then you are supposed to show that \[\large \frac{\frac{1}{b}}{\frac{a}{b}}-\frac{\frac{a}{b}}{\frac{1}{b}}=\frac{b^2}{a}\]

OpenStudy (anonymous):

can you do that algebra?

OpenStudy (anonymous):

yes, but i;m having difficulty coming up with the same denoms

OpenStudy (anonymous):

rewrite the left hand sides as \[\frac{b}{ab}-\frac{ab}{b}\]

OpenStudy (anonymous):

common denominator is \(ab\) so you get \[\frac{b}{ab}-\frac{a^2b}{ab}=\frac{b-a^2b}{ab}\]

OpenStudy (anonymous):

factor and cancel, gets \[\frac{b-a^2b}{ab}=\frac{b(1-a^2}{ab}=\frac{1-a^2}{a}\]

OpenStudy (anonymous):

oh so there is one tiny piece of trig now we have \[\frac{1-\cos^2(x)}{\cos(x)}\] one tiny trig step and you are done